To provide change for a day's business, a merchant buys $75 worth half dollars, quarters, dimes, and nickels. Altogether there are 360 coins, and the number of half dollars is the same as the number of quarters, and the number of dimes is the same as the number of nickels. Find the number of coins of each investments.
Answer by richwmiller(17219)

50*h+25*q+10*d+5*n=7500
h+q+d+n=360
h=q
d=n
50*q+25*q+10*d+5*d=7500
q+q+d+d=360
75*q+15*d=7500
2q+2d=360
q+d=180
q=180-d
5*q+1*d=500
5*(180-d)+1*d=500
900-5d+d=500
400=4d
d=100
n=100
q=80
h=80